Jesse flies a plane against a headwind for 3224 miles. The return trip with the wind took 10 hours less time. If the wind speed is 5 mph, how fast does Jesse fly the plane when there is no wind?
Answer by mananth(16946)

Jesse flies a plane against a headwind for 3224 miles. The return trip with the wind took 10 hours less time. If the wind speed is 5 mph, how fast does Jesse fly the plane when there is no wind?
let speed of plane be x mph
wind speed = 5 mph
against wind time= 3224/(x-4)
with wind time = 3224/(x+4)


3224/(x-4)- 3224/(x+4)=10
simplify
3224(x+4) - 3224(x-4) = 10(x+4)(x-4)
3224x +3224*4 -3224x +3224*4= 10x^2 -160
2*(3224(4)+160 =10x^2
solve for x
x= 50.9 mph speed of plane in still air